OpenStudy (anonymous):

Evaluate each polynomial for x = 3 and for x = −3. 4x2 − 6x + 9

OpenStudy (aroub):

wait.. which one 3 or -3 ??

OpenStudy (anonymous):

replace x by 3 and see what you get i.e. compute \[4\times 3^2-6\times 3+9\] for the first one and \[4\times (-3)^2-6\times (-3)+9\] for the second